# Micro-Diffusion Compression: Binary Tree Tweedie Denoising for Online Probability Estimation

### 09 Mar 2026 v1

@@ Abstract
On the enwik8 benchmark (100 MB Wikipedia), Midicoth achieves 1.753 bpb (21.9 MB), outperforming xz -9 (1.989 bpb) by 11.9% and all evaluated dictionary-based compressors, without any pre-trained neural network, training data, or GPU. On alice29.txt (152 KB Canterbury Corpus), Midicoth achieves 2.119 bpb, outperforming xz -9 (2.551 bpb) by 16.9%. On a 334 KB out-of-distribution government report, Midicoth achieves 1.525 bpb, outperforming xz -9 (1.739 bpb) by 12.3% with a pure online approach.
@@ 4.3 Benchmarks
  3. 3.
    English Indices of Deprivation 2025 (333,794 bytes): Out-of-distribution government report [19], published October 2025—definitively absent from any model’s training data, used to evaluate generalization to unseen text styles.
@@ 5.1 Compression Results
5.2 Out-of-Distribution Evaluation
Table 3: Out-of-distribution evaluation on the English Indices of Deprivation 2025 report (333,794 bytes), published after all models’ training cutoffs.
Compressor Size (B) bpb
Original 333,794 8.000
gzip -9 91,348 2.189
zstd -19 79,709 1.910
xz -9 72,552 1.739
bzip2 -9 69,305 1.661
Brotli -q 11 68,681 1.646
Midicoth (ours) 63,630 1.525
CMIX v21 47,897 1.148
Nacrith 30,171 0.723
Table 3 evaluates on a 334 KB government report published in October 2025. Midicoth achieves 1.525 bpb, outperforming xz -9 (1.739 bpb) by 12.3%, Brotli (1.646 bpb) by 7.4%, and bzip2 (1.661 bpb) by 8.2%—a striking result for a purely online adaptive compressor on an out-of-distribution file.
The ablation on this file (Table 5) reveals that Midicoth’s pipeline is particularly effective on structured prose: the match model contributes +13.1%, and the post-blend Tweedie denoiser adds +2.3%. The full pipeline achieves a +16.6% total improvement over base PPM.
@@ 5.3 Ablation Study
Table 5: Ablation study on English Indices of Deprivation 2025 (333,794 bytes).
Config Size Ratio $\Delta$ Time
Base PPM 76,299 22.86% 0.4s
+ Match 66,342 19.88% +13.05% 0.6s
+ M+Word 65,962 19.76% +0.57% 0.7s
+ M+W+HCtx 65,114 19.51% +1.29% 1.0s
+ M+W+H+Tweedie 63,618 19.06% +2.30% 5.6s
Total impr. +16.62%
@@
Tables 4, 5, and 6 present the incremental contribution of each pipeline layer across three datasets. Key observations:
@@
The match model contributes 5.6% on enwik8_3M and 13.1% on EID, making it the single largest post-PPM component on larger files and highly repetitive data. Its value scales with input repetitiveness and file size.
@@
Applied as the last layer after all model blending, the Tweedie denoiser consistently contributes 2.3–2.8% across all three datasets. This is notably larger than the 1.2–1.6% achieved when Tweedie was applied before blending (i.e., directly to the PPM output). The improvement arises because the blending layers (match, word, high-order context) introduce their own systematic biases—e.g., overconfident match predictions on false matches, or underweighted word predictions on short words—that Tweedie’s calibration tables can detect and correct. By operating on the fully blended distribution, Tweedie calibrates the entire ensemble, not just the PPM model.
@@ 5.4 Cross-Dataset Component Summary
Component alice29 EID enwik3M Avg.
Match +0.57% +13.05% +5.56% +6.39%
Word +1.06% +0.57% +0.75% +0.79%
HighCtx +1.50% +1.29% +2.49% +1.76%
Tweedie +2.63% +2.30% +2.77% +2.57%
Total +5.65% +16.62% +11.14% +11.14%
Table 7 summarizes the cross-dataset contributions. With PPMC exclusion providing a strong base, the post-PPM layers collectively contribute 5–17% additional improvement. The match model shows the highest variance across datasets, reflecting its dependence on input repetitiveness. The post-blend Tweedie layer is notably the most consistent contributor, adding 2.3–2.7% across all three datasets regardless of file type or size.
@@ 7. Conclusion
Third, the full five-layer pipeline achieves 1.753 bpb on enwik8 (100 MB), outperforming xz -9 (1.989 bpb) by 11.9%, 2.119 bpb on alice29.txt, outperforming xz -9 (2.551 bpb) by 16.9%, and 1.525 bpb on the EID 2025 out-of-distribution report, outperforming xz -9 (1.739 bpb) by 12.3%—using only adaptive statistical models with no neural network, no training data, no GPU. This demonstrates that significant compression improvements remain achievable through algorithmic innovation in classical statistical modeling.

### 11 Mar 2026 v2

@@ Abstract
On the enwik8 benchmark (100 MB Wikipedia), Midicoth achieves 1.753 bpb (21.9 MB), outperforming xz -9 (1.989 bpb) by 11.9% and all evaluated dictionary-based compressors, without any pre-trained neural network, training data, or GPU. On alice29.txt (152 KB Canterbury Corpus), Midicoth achieves 2.119 bpb, outperforming xz -9 (2.551 bpb) by 16.9%.
@@ 5.2 Ablation Study
Tables 3 and 4 present the incremental contribution of each pipeline layer across two datasets. Key observations:
@@
The match model contributes 5.6% on enwik8_3M, making it the single largest post-PPM component on larger files and repetitive data. Its value scales with input repetitiveness and file size.
@@
Applied as the last layer after all model blending, the Tweedie denoiser consistently contributes 2.6–2.8% across both datasets. This is notably larger than the 1.2–1.6% achieved when Tweedie was applied before blending (i.e., directly to the PPM output). The improvement arises because the blending layers (match, word, high-order context) introduce their own systematic biases—e.g., overconfident match predictions on false matches, or underweighted word predictions on short words—that Tweedie’s calibration tables can detect and correct. By operating on the fully blended distribution, Tweedie calibrates the entire ensemble, not just the PPM model.
@@ 5.3 Cross-Dataset Component Summary
Component alice29 enwik3M Avg.
Match +0.57% +5.56% +3.07%
Word +1.06% +0.75% +0.91%
HighCtx +1.50% +2.49% +2.00%
Tweedie +2.63% +2.77% +2.70%
Total +5.65% +11.14% +8.40%
Table 5 summarizes the cross-dataset contributions. With PPMC exclusion providing a strong base, the post-PPM layers collectively contribute 5–11% additional improvement. The match model shows the highest variance across datasets, reflecting its dependence on input repetitiveness. The post-blend Tweedie layer is notably the most consistent contributor, adding 2.6–2.8% across both datasets regardless of file type or size.
@@ 7. Conclusion
Third, the full five-layer pipeline achieves 1.753 bpb on enwik8 (100 MB), outperforming xz -9 (1.989 bpb) by 11.9%, and 2.119 bpb on alice29.txt, outperforming xz -9 (2.551 bpb) by 16.9%—using only adaptive statistical models with no neural network, no training data, no GPU. This demonstrates that significant compression improvements remain achievable through algorithmic innovation in classical statistical modeling.